# Alignment-Aware Model Extraction Attacks on Large Language Models

### v1

@@ 4.1 Overview
As described in Algorithm 1, LoRD follows a reinforcement learning paradigm, that is, it consists of several periods, and in each period, the model will learn to explore new responses and attempt to enhance the model trained in the last period. However, different from LLMs’ alignments, the agent can neither obtain the reward from the reward model directly, nor label positive and negative responses manually. This motivates us to develop a locality-reinforced distillation method to address this issue.
Illustrated by Figure 2, the model samples two sentences randomly at period $t-1$, which are denoted as $\mathbf{y}_{t-1}^{+}$ and $\mathbf{y}_{t-1}^{-}$, respectively. In a new period $t$, we first compute the changes of likelihoods for these two sentences, among the old model $P_{\theta_{t-1}}$ and the current model $P_{\theta_{t}}$. These changes of likelihoods, denoted as $\Delta_{t}^{+}$ and $\Delta_{t}^{-}$, indicate whether a selected sentence is locally isotropic ($\Delta>0$) to the victim model’s response $\mathbf{y}_{vic}$ or not ($\Delta\leq 0$), which can be seen as the feedback signal for $P_{\theta_{t}}$ in the current optimization step. For convenience, we may swap $\mathbf{y}_{t-1}^{+}$ with $\mathbf{y}_{t-1}^{-}$ to make sure that $\Delta_{t}^{+}>\Delta_{t}^{-}$ always holds. In this way, for pairs $(\mathbf{x},\mathbf{y}_{vic})$ we can take $\mathbf{y}_{t-1}^{+}$ as a locality neighborhood of $\mathbf{y}_{vic}$ and $\mathbf{y}_{t-1}^{-}$ as the negative sample, all of which can be utilized in the training of $P_{\theta_{t}}$. Figure 3 illustrates this procedure. Additionally, we take $\mathbf{y}_{t-1}^{+}$ as the positive label under the current scope only when $\Delta^{+}$ and $P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})$ both exceed their respective fixed thresholds $\tau{1}$ and $\tau_{2}$. If these conditions are not met, we will use $\mathbf{y}_{vic}$ as a substitute for $\mathbf{y}_{t-1}^{+}$ to enable a cold start.
Based on $\mathbf{y}_{vic}$, $\mathbf{y}_{t-1}^{+}$, and $\mathbf{y}_{t-1}^{-}$, we now derive the loss function of LoRD.
4.2 Derivation of Loss Functions
@@
Objective function $\mathcal{L}_{obj}$. Following the definition of $R_{\theta_{\phi}}$ in Equation 6, we use the logarithmic proportion between the positive sample and the negative sample as the de-biased reward, i.e., $\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$. Subsequently, we take $\text{log}[P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})/P_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x})]$, the proportion of the victim model to the local model, as the ratio term to scale such a reward.
Therefore, the objective function can be formatted as
¯$\bar{\mathcal{L}}_{obj}$ $=-\sum_{\mathbf{x}\in\mathcal{D}_{q}}|\text{log}[\frac{P_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})}]|\cdot\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}]$ (8)
$=-\sum_{\mathbf{x}\in\mathcal{D}_{q}}[\frac{\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{|\text{log}P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})-\text{log}P_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x})|}],$
where ||$\cdot$|| ⋅ | indicates no gradient back propagation, $t$ and $\mathbf{x}\sim\mathcal{D}_{q}$ denote the training period and the sampled query text, respectively. Since it is impractical to obtain the victim model’s probability $P_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x})$ under black-box settings, we simplify Equation 8 by setting $P_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x})=1$, and remove $P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})$ to ensure the stability in the black-box extraction, i.e.,
@@
Regularization loss $\mathcal{L}_{reg}$. Current studies [44, 41, 7] tend to constrain $\theta_{t}$ with initial model’s generating distribution $P_{\theta_{init}}(\cdot|\mathbf{x})$ in RLHF. However, this strategy requires re-loading an extra initial model’s weights, which is memory-consuming. Also, the KL-divergence term $\mathbb{D}_{KL}(P_{\theta_{t}}||P_{\theta_{init}})$ consists of $P_{\theta_{t}}(\cdot|\mathbf{x})$, which requires an extra exponential operation of $\text{log}P_{\theta_{t}}(\cdot|\mathbf{x})$ that will slow down the forward computation procedure.111logsoftmax is preferred in the implementation of deep learning frameworks [6], as the exponential operation in softmax and the logarithmic operation in cross-entropy can be canceled out by each other. Therefore, in MEAs, we put forward a regularization based on the victim model, where only the logarithmic part of KL divergence is employed and only the dimension of selected tokens is used in KL divergence, i.e.,
@@
In Equation 10, we utilize a $clip(\cdot)$ function to limit the value of the regularization term, as we expect the regularization term could only be used to avoid the off the cliff problem [45, 44] in RL’s convergence and the shift of the distribution [41] in LLMs’ alignments.
@@
Finally, we wrap $\mathcal{L}_{\text{LoRD}}$ with a sigmoid function $\sigma(\cdot)$ to normalize the loss to the interval $(0,1)$, which is
$\mathcal{L}=\mathbb{E}~{}\sigma(\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})}]+clip(\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{{vic}}|\mathbf{x})}])).$ (12)
@@ 5 Theoretical Analysis
This section will compare LoRD with current model extraction methods from a theoretical perspective. We will first reveal the underlying inconsistency between the optimization procedure of LLMs, which typically involves RL-based alignments, and the previous model extraction approaches utilizing MLE and knowledge distillation (KD). Subsequently, we will demonstrate in theory the reasons why LoRD can achieve stronger watermark resistance and higher query efficiency than existing methods.
5.1 Consistency Analysis regarding Different Learning Tasks
Figure 4: Illustrations for the converging procedure of probability distributions regarding four different methods, namely MLE (a), KD (b), RLHF (c), and LoRD (d). Arrows indicate the expected optimization direction. We mark the distribution dimensions learned with labels in blue, and employ pink and yellow components to indicate the probabilities of positive and negative tokens, respectively. Dimensions without any explicit constraints in the loss function are in $white$, though value changes can occur during training.
[…]
Expected Distribution of KD. Following a previous work [21], the objective function of KD is
where $\text{SM}(\cdot)$ represents the softmax function, and $T>1$ denotes the temperature to smooth the targeted distribution $P_{\theta_{vic}}(\cdot|\mathbf{x})$. As described in Equation 14, knowledge distillation aims to align $P_{\theta}(\cdot|\mathbf{x})$ with $P_{\theta_{vic}}(\cdot|\mathbf{x})$ in both the original and the smoothed probability across all dimensions, which is exceptionally comprehensive among these methods.
where $\theta*$ denotes the expected parameters of the models as
More detailed derivation is in Appendix C.1. Replacing Equation 15 with Equation 16, the expected distribution can be represented as $\mathbf{r}_{i,j}\cdot P_{\theta_{init}}(\cdot|\mathbf{x})$, in which $\mathbf{r}_{i,j}$ indicates the wrapped distribution gain. This distortion aims to maximize the ratio $P_{\theta}(y^{+}_{j}|\mathbf{x},\mathbf{y}^{+}_{<j})/P_{\theta}(y^{-}_{j}|\mathbf{x},\mathbf{y}^{-}_{<j})$, and leave the probabilities in other dimensions unconstrained directly.
Consistency Analysis of Four Objective Functions. Based on the analysis of the above four objective functions, we illustrate their convergence procedure, as exhibited in Figure 4. We reach the following proposition.
@@ 5.1 Consistency Analysis regarding Different Learning Tasks
Albeit the inconsistency in their training procedures, the following theorem demonstrates that with enough samples, all these methods will reach the same distribution results.
Theorem 1 (Equivalence when Converged).
Ideally, for any loss value of Equations 4, 5, 6, 11, or 12 converging to 00, we have $\mathbf{y}^{+}\equiv\mathbf{y}_{vic}$. Meanwhile, the local model’s distribution $P_{\theta}(\cdot|\mathbf{x})$ will approach that of the victim model $P_{\theta_{vic}}(\cdot|\mathbf{x})$ on MEAs from all three discussed MEA methods, including LoRD, MLE, and KD.
Theorem 1 ensures that the local model will converge to the victim model regardless of the choice of MEA methods. Nonetheless, by Proposition 1, LoRD still has two benefits: query time reduction (i.e., size of the query set $\mathcal{D}_{q}$), and the watermark resistance of the learned local model. We will elaborate on them in Section 5.2.
@@ 5.2 Comparative Analysis on Model Stealing
Query Efficiency. Let $N_{Q}$ and $N_{R}$ denote the sequence lengths of the query text and the response text, respectively. For MLE, the ideal query numbers to populate the entire text space are given by $\mathcal{O}(V^{N_{Q}}\cdot V^{N_{R}})$, where $V$ represents the size of the vocabulary. In contrast, the ideal query times for LoRD can be reduced to $\mathcal{O}(V^{N_{Q}}\cdot C)$, with $C$ being a constant. This constant suggests that the model requires only a fixed and limited number per query, as LoRD itself can automatically generate and explore $V^{N_{R}}$ possibilities during training.
Table 1: Datasets and pre-trained model checkpoints used in the paper.
  BLEU BERTScore Rouge-L
1 2 3 4 Pre. Rec. F1. Pre. Rec. F1.
Text to SQL: WikiSQL [58] with 64 query samples
Victim Model 54.1 41.4 32.1 24.4 86.9 93.5 90.1 58.9 62.1 59.7
Text to SQL: Spider [58] with 64 query samples
Victim Model 9.4 3.9 2.1 1.1 77.7 84.1 80.6 17.1 36.3 21.8
Data to Text: E2E NLG [14] with 64 query samples
Victim Model 51.8 27.0 26.8 19.1 93.9 94.6 94.2 49.6 54.6 51.4
Data to Text: CommonGen [31] with 64 query samples
Victim Model 33.3 18.5 11.1 6.9 91.3 92.1 91.7 33.6 40.7 36.1
Summarization: TLDR [27] with 64 query samples
Victim Model 11.9 5.0 2.6 1.5 85.9 88.4 87.1 13.4 30.9 18.4
Summarization: CNN Daily Mail [20] with 64 query samples
Victim Model 20.4 10.8 6.4 4.1 86.4 87.8 87.1 22.4 40.8 28.2
Summarization: Samsum [17] with 64 query samples
Victim Model 20.7 11.4 6.9 4.4 88.1 91.7 89.8 24.2 50.5 31.6
Table 2: MEA comparison on three tasks, including structured text generation, data to text, and summarization. We use GPT-3.5-turbo as the victim model, and Llama3-8B [1] as the local initial model.
Based on the above analysis, a straightforward concern with employing MLE in LLMs’ extraction is that, given the limited query times in real-world practices, it may suffer from incomplete learning, especially for text generation tasks. Consequently, the local model may tend to memorize some specific responses instead of achieving a broad understanding and generation. We call such a phenomenon preference overfitting (PO), which indicates that the local model is only effective on a limited set of explored samples, and yet does not generalize well to unseen scenarios. In such cases, the local model usually exhibits a more “rugged” decision surface, which appears to overfit the preference sentences in $\mathcal{D}_{tr}$, as shown in Figure 5 (b). Figure 11 provides the empirical visualization of this analysis.
Watermark Resistance. Another limitation of prevalent objective functions, such as MLE and KD, is their susceptibility to watermarks [12, 19, 18, 25] of output contents, i.e., while stealing knowledge from LLMs via responses $\mathbf{y}_{vic}$, watermarks within them will also been passively inherited by the local model. Consequently, the generated sentences of the local model may possess some residual of watermarks, which might be detected as evidence of stealing.
LoRD can mitigate the influences of watermarks naturally, as it does not learn the likelihood of victim models’ responses $\mathbf{y}_{vic}\sim\mathcal{D}_{tr}$ directly, but relies on $\mathbf{y}_{vic}$ to determine positive and negative labels from responses generated by the local model.
@@
$\mathcal{L}=$ $\mathbb{E}[(1-\lambda_{1})\cdot(\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x}))$ (17)
$\quad+\lambda_{1}\cdot clip(\text{log}P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})),$
@@
When $\lambda_{1}$ is small, the convergence of LoRD will substantially focus on maximizing $P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})/P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$, with which the local model will exhibit a strong watermark resistance ability. When $\lambda_{1}$ increases, LoRD will tend to rely more on the guidance of $\mathbf{y}_{vic}$, resulting in a higher risk of introducing watermarks. In the case of $\lambda_{1}=1$, the local model will converge to the victim model without any exploration and watermark resistance, which might suffer from the defense by watermarks.
From a global perspective, $\mathcal{L}_{obj}$ represents the exploration and the locality learning ability of LoRD, which can mitigate the influences of watermarks. On the other hand, $\mathcal{L}_{reg}$ ensures the stability of the training procedure. Therefore, Equation 17 characterizes a trade-off via $\lambda_{1}$ between the stability and the diversity during stealing, and Equation 12 can be seen as a special case of Equation 17 with $\lambda_{1}=0.5$.

### v2

@@ 3.1 Overview
Figure 3: Determination of the positive and negative samples in LoRD. We sample $\mathbf{y}_{t-1}^{+}$ and $\mathbf{y}_{t-1}^{-}$ from $P_{\theta_{t-1}}(\cdot|\mathbf{x})$, and compute their conditional probabilities. The response with a higher probability increment on $\theta_{t}$ is selected as the positive sample.
In this subsection, we delve into the details of our model extraction framework, LoRD (Locality Reinforced Distillation). As described in Algorithm 1, LoRD follows a reinforcement learning paradigm, that is, it consists of several periods, and in each period, the model will learn to explore new responses and attempt to enhance the model trained in the last period. However, different from LLMs’ alignments, the agent can neither obtain the reward from the reward model directly, nor label positive and negative responses manually. This motivates us to design a new RL method which can implicitly measure the reward for generated tokens under the guidance of victim model’s responses.
Illustrated by Figure 2, LoRD first requires the model to sample two sentences randomly at period $t-1$, which are denoted as $\mathbf{y}_{t-1}^{+}$ and $\mathbf{y}_{t-1}^{-}$, respectively. In a new period $t$, it first computes the changes of likelihoods for these two sentences, among the old model $P_{\theta_{t-1}}$ and the current model $P_{\theta_{t}}$. These changes of likelihoods, denoted as $\Delta_{t}^{+}$ and $\Delta_{t}^{-}$, indicate whether a selected sentence is locally isotropic ($\Delta>0$) to the optimization direction with victim model’s response $\mathbf{y}_{vic}$ or not ($\Delta\leq 0$), which can be seen as the feedback signal for $P_{\theta_{t}}$ in the current optimization step. For convenience, we may swap $\mathbf{y}_{t-1}^{+}$ with $\mathbf{y}_{t-1}^{-}$ to make sure that $\Delta_{t}^{+}>\Delta_{t}^{-}$ always holds. In this way, for pairs $(\mathbf{x},\mathbf{y}_{vic})$ we can take $\mathbf{y}_{t-1}^{+}$ as a locality neighborhood of $\mathbf{y}_{vic}$ and $\mathbf{y}_{t-1}^{-}$ as the negative sample, all of which can be utilized in the training of $P_{\theta_{t}}$. Figure 3 illustrates this procedure. Additionally, LoRD takes $\mathbf{y}_{t-1}^{+}$ as the positive label under the current scope only when $\Delta^{+}$ or $P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})$ exceed their respective fixed thresholds $\tau_{1}$ and $\tau_{2}$. If these conditions are not met, it will use $\mathbf{y}_{vic}$ as a substitute for $\mathbf{y}_{t-1}^{+}$ to enable a cold start.
Based on $\mathbf{y}_{vic}$, $\mathbf{y}_{t-1}^{+}$, and $\mathbf{y}_{t-1}^{-}$, we now design LoRD’s loss function.
3.2 Design of Loss Functions
@@
Objective function $\mathcal{L}_{obj}$. Inspired by the reward model $R_{\theta_{\phi}}$ existed in Equation 6, which is trained to distinguish between positive and negative samples, we propose utilizing the logarithmic proportion of positive to negative samples as the means of achieving a de-biased reward, i.e.,
@@
Equation 8 exhibits similarities to previous studies on RL-enhanced LLM (Peters and Schaal, 2007; Peng et al., 2019; Go et al., 2023; Korbak et al., 2022; Rafailov et al., 2023). We provide a theoretical explanation for its consistency with the learning procedure of RLHF and the deduction procedure, as detailed in Section 4 and Appendix B.1.
However, training the local model merely by $\mathcal{L}_{obj}$ is ineffective due to two reasons: i) when  $\mathcal{L}_{\text{ LoRD}}:=\mathcal{L}_{obj}$, no information from the victim model’s responses is incorporated into the selection of $\mathbf{y}^{+}_{t-1}$ beyond the cold start phase, resulting in a meaningless self-reward-based learning loop for the stealing procedure; ii) the convergence of the local model’s training cannot be guaranteed.
To address these two issues simultaneously, we design the regularization term as follows.
Regularization loss $\mathcal{L}_{reg}$. Different from LLM’s RLHF (Schulman et al., 2015; Rafailov et al., 2023; Bai et al., 2022a) that typically constrain $\theta_{t}$ with initial model’s generating distribution $P_{\theta_{init}}(\cdot|\mathbf{x})$, LoRD aims to directly constrain $\theta_{t}$ with victim model’s distribution $P_{\theta_{vic}}(\cdot|\mathbf{x})$.
Unfortunately, $P_{\theta_{vic}}(\cdot|\mathbf{x})$ is typically inaccessible within the APIs of commercial LLMs and is not feasible for our black-box scenarios. Consequently, we incorporate the regularization techniques employed in PPO and TRPO but tailor our regularization as a bounded contrastive term between the likelihood of $\theta_{t}$ under the victim model’s response and the negative sample, i.e.,
@@
In Equation 9, we utilize PPO’s $clip(\cdot)$ function to limit the value of the regularization term, as we expect the regularization term could only be used to avoid the off the cliff problem (Schulman et al., 2017, 2015) in RL’s convergence. Besides, our contrastive term can be seen as a streamlined black-box variant of the KL divergence in TRPO. This simplification offers two advantages: i) it alleviates the necessity of loading the initial model’s weights, leading to a substantial reduction in GPU memory usage; ii) it eliminates the need for $P_{\theta_{t}}(\cdot|\mathbf{x})$, which would otherwise necessitate an additional exponential operation of $\text{log}P_{\theta_{t}}(\cdot|\mathbf{x})$ that slow down the forward process and increase extra consumption.111logsoftmax is preferred in the implementation of deep learning frameworks (PyTorch, 2024), as the exponential operation in softmax and the logarithmic operation in cross-entropy can be canceled out by each other.
@@
Finally, we wrap $\mathcal{L}_{\text{LoRD}}$ with a sigmoid function $\sigma(\cdot)$ to normalize the loss to the interval $(0,1)$:
$\mathcal{L}=\sum_{\mathbf{x}\sim\mathcal{D}_{q}}~{}\sigma(\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})}]+clip(\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{{vic}}|\mathbf{x})}])).$ (11)
@@ 4 Theoretical Analysis
This section will compare LoRD with current model extraction methods from a theoretical perspective. We will first reveal the underlying inconsistency between the optimization of LLMs, which typically involves RL-based alignments, and the previous MEAs utilizing MLE and KD. Subsequently, we will demonstrate the reasons why LoRD can achieve stronger watermark resistance and higher query efficiency than existing methods.
4.1 Consistency Analysis on Learning Tasks
Based on the analysis of the four objective functions for MLE, KD, RLHF and LoRD, we reach the Proposition 1, and illustrate their convergence procedure exhibited in Figure 4. A detailed explanation for it can be found in Appendix B.1.
Figure 4: Illustrations for the converging procedure of probability distributions regarding four methods, namely MLE (a), KD (b), RLHF (c), and LoRD (d). Arrows indicate the expected optimization direction. We mark the distribution dimensions learned with labels in blue, and employ pink and yellow components to indicate the probabilities of positive and negative tokens, respectively.
@@ 4.1 Consistency Analysis on Learning Tasks
Albeit the inconsistency in their training procedures, we put forward Proposition 2 to demonstrate that with enough samples, all these methods will reach the same distribution results.
Proposition 2 (Equivalence when Converged).
Ideally, for any loss value of Equations 4, 5, 6, 10, or 11 converging to 00, we have $\mathbf{y}^{+}\equiv\mathbf{y}_{vic}$. Meanwhile, the local model’s distribution $P_{\theta}(\cdot|\mathbf{x})$ will approach that of the victim model $P_{\theta_{vic}}(\cdot|\mathbf{x})$ on MEAs from all three discussed MEA methods, including LoRD, MLE, and KD.
Proposition 2 ensures that the local model will converge to the victim model regardless of the choice of MEA methods. So what is the benefit of LoRD? In Section 4.2, we will show that LoRD outperforms current MEAs with two aspects: the query time reduction, and the watermark resistance of the learned local model.
@@ 4.2 Comparative Analysis on Model Stealing
Query Efficiency. Let $N_{Q}$ and $N_{R}$ denote the sequence lengths of the query text and the response text, respectively. For MLE, the ideal query numbers to populate the entire text space are given by $\mathcal{O}(V^{N_{Q}}\cdot V^{N_{R}})$, where $V$ represents the size of the vocabulary. In contrast, LoRD possesses the capability to automatically explore the generation token space, thereby significantly reducing the query requirements about generation candidates to a constant level. Specifically, the complexity of LoRD’s query requirements is $\mathcal{O}(V^{N_{Q}}\cdot C)$, where $C$ is a constant that correlates with the capability of local models.
Based on the above analysis, a straightforward concern with employing MLE in LLMs’ extraction is that, given the limited query times in real-world practices, it may suffer from incomplete learning, especially for text generation tasks. Consequently, the local model may tend to memorize some specific responses instead of achieving a broad understanding and generation. We call such a phenomenon preference overfitting (PO), which indicates that the local model is only effective on a limited set of explored samples, and yet does not generalize well to unseen scenarios. In such cases, the local model usually exhibits a more “rugged” decision surface, which appears to overfit the preference sentences in $\mathcal{D}_{tr}$, as shown in Figure 11 (b). Figure 10 provides a visualization of it.
Watermark Resistance. Another limitation of prevalent objective functions, such as MLE and KD, is their susceptibility to watermarks (Cong et al., 2022; He et al., 2022, 2021; Kirchenbauer et al., 2023) of output contents, i.e., while stealing knowledge from LLMs via responses $\mathbf{y}_{vic}$, watermarks within them will also been passively inherited by the local model. Consequently, the generated sentences of the local model may possess some residual of watermarks, which might be detected as evidence of stealing.
Despite introducing current watermark removal techniques, we indicate that LoRD can mitigate the influences of watermarks naturally, as it does not learn the likelihood of victim models’ responses $\mathbf{y}_{vic}\sim\mathcal{D}_{tr}$ directly, but relies on $\mathbf{y}_{vic}$ to determine positive and negative labels from responses generated by the local model.
@@
$\mathcal{L}$ $=\mathbb{E}[(1-\lambda_{1})\cdot(\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x}))$
$\quad\quad~{}~{}+\lambda_{1}\cdot clip(\text{log}P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x}))],$
@@
When $\lambda_{1}$ is small, the convergence of LoRD will substantially focus on maximizing $P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})/P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$, with which the local model will exhibit a strong watermark resistance ability. When $\lambda_{1}$ increases, LoRD will tend to rely more on the guidance of $\mathbf{y}_{vic}$, resulting in a higher risk of introducing watermarks. In the case of $\lambda_{1}=1$, the local model will converge to the victim model without any exploration and watermark resistance, which might suffer from the same level of defense by watermarks.
From a global perspective, $\mathcal{L}_{obj}$ represents the exploration and the locality learning ability of LoRD, which can mitigate the influences of watermarks. On the other hand, $\mathcal{L}_{reg}$ ensures the stability of the training procedure. Therefore, $\mathcal{L}$ characterizes a trade-off via $\lambda_{1}$ between the stability and the diversity during stealing, and Equation 11 can be seen as a special case of $\mathcal{L}$ with $\lambda_{1}=0.5$.
We provide an empirical comparison for query efficiency in Appendix A.2.1, and the comparison on watermark resistance in Appendix A.1.